# Understanding Addition In Transformers

### 29 Feb 2024 v6

@@ Abstract
Understanding the inner workings of machine learning models like Transformers is vital for their safe and ethical use. This paper presents an in-depth analysis of a one-layer Transformer model trained for $n$-digit integer addition. We reveal that the model divides the task into parallel, digit-specific streams and employs distinct algorithms for different digit positions. Our study also finds that the model starts calculations late but executes them rapidly. A rare use case with high loss is identified and explained. Overall the model’s algorithm is explained in detail. These findings are validated through rigorous testing and mathematical modeling, contributing to the broader works in Mechanistic Interpretability, AI safety, and alignment. Our approach [^1] To encourage re-use and reproducibility we made out code available here:
https://github.com/apartresearch/Interger_Addition
opens the door for analyzing more complex tasks and multi-layer Transformer models.
@@ 1 Introduction
Understanding the underlying mechanisms of machine learning models is essential for ensuring their safety and reliability [Barez et al., 2023, Olah et al., 2020b]. Specifically, the sub-field of mechanistic interpretability within machine learning interpretability aims to dissect the behavior of individual neurons and their interconnections in neural networks [Räuker et al., 2022]. This pursuit is part of a larger endeavor to make the decision-making processes of complex machine learning models transparent and understandable. Although models like Transformers have shown remarkable performance on a myriad of tasks, their complexity makes them challenging to interpret. Their multi-layered architecture and numerous parameters make it difficult to comprehend how they derive specific outputs [Vig, 2019]. Further, while simple arithmetic tasks like integer addition may be trivial for humans, understanding how a machine learning model like a Transformer performs such an operation is far from straightforward.
@@
Our results demonstrate the model’s unique approach applies to integer addition across various digit lengths (Refer Appendixes B and C). Our theoretical framework provides a mathematical justification for the model’s behavior, substantiating our empirical observations and offering a foundation for future work.
Our main contributions are:
@@
Figure 1: Illustration of the transformer model’s attention pattern when adding two 5-digit integers. The model attends to digit pairs sequentially from left to right, resulting in a "double staircase" pattern across rows. A: The 5 digit question is revealed token by token. The “10s of thousands" digit is revealed first. B: From the “=" token, the model attention heads focus on successive pairs of digits, giving a “double staircase" attention pattern. C: The 3 heads are time-offset from each other by 1 token such that, in each row, data from 3 tokens is available. D: To calculate A3, the 3 heads do independent simple calculations on D3, D2 and D1. The results are combined by the MLP layer using trigrams. A3 is calculated one token before it is needed. This approach applies to all answer digits, with the first and last digits using slight variations of the approach.
@@ 2 Background
We focus on a single-layer transformer model with a vocabulary of size $V$ containing the symbols “0” to “9", “+" and “=". The model converts the human readable input (e.g. “12345+67890=") into an input sequence $(\texttt{x}_{1},\ldots,\texttt{x}_{p})$ where each $\texttt{x}_{i}\in\{1,\ldots,V\}$. Tokens are mapped to de dimensional embeddings by selecting the $\texttt{x}_{i}$-th column of $E\in\mathbb{R}^{d_{e}\times V}$. The model processes the input tokens, using a mechanism called “self-attention". Each input token is passed through a self-attention mechanism that calculates weighted relationships between all input tokens - capturing the importance of each token relative to others. The model then aggregates these weighted representations to produce contextually enriched representations for each token. These enriched representations are subsequently fed through feedforward neural networks (i.e. an MLP) to refine their information. Finally, the output tokens are generated based on the refined representations, and converted back to human readable format using the vocabulary (e.g. “12345+67890=80235").
@@ 3 Related Work
Interpreting and reverse engineering neural networks and transformers to find meaningful circuits has been an area of active research. Olah et al. [2020a] argued that by studying the connections between neurons and their weights, we can find meaningful algorithms (aka Circuits) in a “vision” neural network. Elhage et al. [2021] extended this approach to transformers, conceptualizing their operation in a mathematical framework that allows significant understanding of how transformer operate internally. Various tools [Foote et al., 2023, Conmy et al., 2023b] use this framework to semi-automate some aspects of reverse engineering. Nanda et al. [2023] reverse-engineered modular addition (e.g. 5 + 7 mod 10 = 2) showing the model used discrete Fourier transforms and trigonometric identities to convert modular addition to rotation about a circle.
@@
Figure 2: Training loss curves per digit position for a 5-digit integer addition task, showing the model trains each digit semi-independently.
Circuit analysis can extract graphical circuit representations and analyze component interactions [Bau et al., 2017]. To enable analysis and interpretability, techniques in works like Petersen et al. [2021] symbolically reverse-engineer networks by recovering computational expressions. Research including Seth [2005] advocate analyzing networks causally, introducing structural causal models to infer mechanisms. Examinations of sequence models like Petersen et al. [2021] have analyzed the emergence and interaction of modular components during training. Evolutionary perspectives such as Miikkulainen [2021] elucidate how selection pressures shape hierarchical representations.
@@ 4 Methodology
The integer addition problem space is very dense. For 5 digit addition, there are 10 billion distinct questions (e.g. “54321+77779=") . The model must predict all 6 answer digits correctly to get the one right answer out of 200,000 possibilities. Changing a single digit in the question changes 1 to 6 digits in the answer. The full question is only revealed one token (the “=") before the model must predict the first answer digit.
Our model was trained on 1.8 million out of 10 billion questions. After training, the model predicts answers to questions with low loss, showing the model does not rely on memorisation of training data.
Fig. 2 shows the model trains each digit semi-independently suggesting the model performs integer addition by breaking down the task into parallel digit-specific streams of computation.
The traditional human addition process first sums the units before moving on to higher value digits. This is the simplest process but relies on being able to choose the order to process the digits in. This autoregressive transformer model processes text from left to right. So the model sees the higher value digits (e.g. thousands) of the question before the lower value digits (e.g. units). It doesn’t use the traditional process.
A key component of addition is the need to sum each digit in the first number with the corresponding digit in the second number. Transformer models contain “attention heads" and they are the only computational sub-component of a model that can move information between positions (aka digits or tokens). Visualising which token(s) each attention head focussed on in each row of the calculation provided insights. While our model works with 2, 3 or 4 attention heads, 3 attention heads gives the most easily interpreted attention patterns. Fig. 3 shows the attention pattern for a single 5 digit addition calculation using 3 attention heads. Appendix C shows the same pattern for 10 and 15 digit addition. Appendix C shows the pattern with 2 or 4 attention heads.
@@ 5 Mathematical Framework
To help investigate, we created a mathematical framework describing what any algorithm must do if it is to perform addition correctly.
Our intuition is that the model a) incrementally discovers a necessary and sufficient set of addition sub-tasks (minimising complexity), b) discovers these sub-tasks semi-independently (maximising parallelism), and c) treats each digit semi-independently (more parallelism). Our framework reflects this.
To explain the framework, let us first define $x$ and $y$ be two $n$-digit integers that need to be added, represented as vectors where $x=(x_{0},x_{1},\ldots,x_{n-1})$ and $y=(y_{0},y_{1},\ldots,y_{n-1})$.
We assert that the framework utilizes three base functions that operate on individual digit pairs. The first is Base Add (aka BA ), which calculates the sum of two digits $x_{i}$ and $y_{i}$ modulo 10, ignoring any carry over from previous columns. The second is Make Carry 1 (aka MC1 ), which evaluates if adding digits $x_{i}$ and $y_{i}$ results in a carry over of 1 to the next column. The third is Make Sum 9 (aka MS9 ), which evaluates if $x_{i}+y_{i}=9$ exactly.
@@
These tasks occur in the training data with different, predictable frequencies (e.g. BA is common, US9 is rarer). Compound tasks are reliant on the base tasks and so discovered later in training. The discovery of each task reduces the model loss by a different, predictable amount (e.g. BA by 50%, US9 by 5%). Combining these facts give an expected order of task discovery during training.
@@
We use this mathematical framework solely for analysis to gain insights. The model training and all loss calculations are completely independent of this mathematical framework.
@@ 6 Training Analysis
Fig. 2 shows the model trains each digit semi-independently. Armed with the mathematical framework, we investigated each digit separately.
The Digit 0 calculation is the least interesting as it only uses BA (not UC1 or US9 ). Once discovered, Digit 0 always quickly refines to have the lowest loss and least noise (as expected). (Graphs in Appendix B.)
For the other digits, we categorised the training data into 3 non-overlapping subsets aligned to the BA , UC1 and US9 tasks, and graphed various combinations, finding interesting results.
The US9 graphs are much noisier than other graphs (Fig. 5). We found that the model has low loss on simple US9 cases (e.g. 45 + 55 = 100) but has high loss on US9 cascades (e.g. 445 + 555 = 1000) where the MC1 must be propagated "right to left" two 2, 3 or 4 columns. The model can’t perform these rare use cases safely, as it has a "left to right" algorithm.
@@
Both the BA and UC1 tasks need to move data between tokens, and so will be implemented in attention head(s). Fig. 6 shows they are trained semi-independently. We choose the number of attention heads in our model with the clearest separation of tasks in the attention pattern. We find (later) that our model has separate attention heads for the BA and UC1 tasks.
Digit 4, the highest question digit, has a significantly different loss curve (shown in Fig. 7) than Digits 1, 2 and 3. This is partially explained by Digit 4 only having simple US9 cases (i.e. no US9 cascades). This does not explain the BA or UC1 differences. This difference persists with different seed values, and with 10 or 15 digit addition. We explain this difference later.
@@ 7 Prediction Analysis
During model prediction, we overrode (mean ablated) the model memory (residual stream) at each row, and confirmed that the addition algorithm does not use any data generated in rows 0 to 10 inclusive. In these rows the model has not yet seen the full question and every digit in the question is independent of every other digit, making accurate answer prediction infeasible. The model also does not use the last (17th) row. Therefore, the addition is started and completed in 6 rows (11 to 16). Further (ablation) experiments confirmed that the A0 to A4 answers are calculated one row before being revealed. (Details in Appendix H.)
The model has slightly different algorithms for the first digit pairs, the middle digit pairs and the last digit pairs. Fig. 1 has a simplified version of how the model calculates the middle digit pair A3. Fig. 8 has more details. For 5 digit addition, there are 2 middle digit pairs (A3 and A2) whereas for 15 digit addition there are 12 middle digit pairs.
Figure 8: A: For A3, the addition algorithm must combine information from digits 3, 2 and 1. B: The 1st head calculates MC1 on digit 1. C: The 2nd head calculates MC1 and MS9 (at most one of which can be true) on digit 2. D: The 3rd head calculates Base Add on digit 3. E: The MLP layer uses trigrams to combine the information from the 3 heads to give the final answer A3, one row before it is output. Appendix G shows this algorithm as pseudocode.
The A3 addition algorithm has three clauses related to digits 3, 2 and 1. Ablating each head in turn shows that the 3rd head has most impact on loss, the 2nd head has less impact, and the 1st head has little impact. This aligns with the intuition that the sum “D3 + D3’" matters most, the MC1 from the previous digit (D2 + D2’) matters less, and the rare MC1 from the previous previous digit (D1 + D1’) matters least.
The last two digits, A1 and A0, use a simplified a version of the A3 algorithm, with some of the three clauses not needed.
The A3 algorithm could also successfully be applied to A4. But the Digit 4 training curve is better (faster) than the middle digits. The attention patterns shows that for A4, the model is using all the heads in row 11 (the “=" token) when the A3 algorithm doesn’t require this. Uniquely, A4 utilises more “compute" than is available to A3, A2, A1 or A0. We assume the model uses this advantage to implement a faster-training and lower-loss algorithm for A5 and A4. We haven’t worked out the details of this.
@@
Despite being feasible, the model does not calculate the task MC1 in rows 7 to 11. Instead it completes each digit calculation in 1 row, possibly because there are training optimisation benefits in generating a “compact" algorithm.
@@
8 Conclusions and Future Work

### 18 Mar 2024 v8

@@ Abstract
Understanding the inner workings of machine learning models like Transformers is vital for their safe and ethical use. This paper provides a comprehensive analysis of a one-layer Transformer model trained to perform $n$-digit integer addition. Our findings suggest that the model dissects the task into parallel streams dedicated to individual digits, employing varied algorithms tailored to different positions within the digits. Furthermore, we identify a rare scenario characterized by high loss, which we explain. By thoroughly elucidating the model’s algorithm, we provide new insights into its functioning. These findings are validated through rigorous testing and mathematical modeling, thereby contributing to the broader fields of model understanding and interpretability. Our approach opens the door for analyzing more complex tasks and multi-layer Transformer models.
footnotetext: Corresponding author: fazl@robots.ox.ac.uk.
@@ 1 Introduction
Understanding the underlying mechanisms of machine learning models is essential for ensuring their safety and reliability [Barez et al., 2023, Olah et al., 2020b, Doshi-Velez and Kim, 2017, Hendrycks and Mazeika, 2022]. By unraveling the inner workings of these models, we can better understand their strengths, limitations, and potential failure modes, enabling us to develop more robust and trustworthy systems. Specifically, the sub-field of mechanistic interpretability within machine learning interpretability aims to dissect the behavior of individual neurons and their interconnections in neural networks [Räuker et al., 2022]. Recent interpretability work has explored how transformers make predictions [Neo et al., 2024], analyzed reward model divergence in large language models [Marks et al., 2024], and highlighted the importance of such analyses for measuring value alignment [Barez and Torr, 2023]. This pursuit is part of a larger endeavor to make the decision-making processes of complex machine learning models transparent and understandable.
Although models like Transformers have shown remarkable performance on a myriad of tasks, their complexity makes them challenging to interpret. Their multi-layered architecture and numerous parameters make it difficult to comprehend how they derive specific outputs [Vig, 2019]. Further, while simple arithmetic tasks like integer addition may be trivial for humans, understanding how a machine learning model like a Transformer performs such an operation is far from straightforward [Liu and Low, 2023].
@@
Our results demonstrate the transformer’s unique approach applies to integer addition across various digit lengths (see Appendixes B and C). This transformer architecture, with its self-attention mechanism and ability to capture long-range dependencies, offers a powerful and flexible framework for modeling sequential data. Our theoretical framework provides a mathematical justification for the model’s behavior, substantiating our empirical observations and offering a foundation for future work in this domain.
Our main contributions are four-fold:
@@
Figure 1: Illustration of the transformer model’s attention pattern when adding two 5-digit integers. The model attends to digit pairs sequentially from left to right, resulting in a “double staircase" pattern across rows. A: The 5 digit question is revealed token by token. The “10s of thousands" digit is revealed first. B: From the “=" token, the model attention heads focus on successive pairs of digits, giving a “double staircase" attention pattern. C: The 3 heads are time-offset from each other by 1 token such that, in each row, data from 3 tokens is available. D: To calculate A3, the 3 heads do independent simple calculations on D3, D2 and D1. The results are combined by the MLP layer using trigrams. A3 is calculated one token before it is needed. This approach applies to all answer digits, with the first and last digits using slight variations of the approach.
@@ 2 Background
We focus on a single-layer transformer model with a vocabulary of size $V$ containing a set of symbols $\mathcal{V}$. The model converts input (e.g., a sequence of symbols from $\mathcal{V}$) into an input sequence $(\mathbf{x}_{1},\ldots,\mathbf{x}_{p})$, where each $\mathbf{x}_{i}\in\mathbb{R}^{V}$ is a one-hot vector representing the corresponding symbol from $\mathcal{V}$. The input tokens are mapped to $d_{e}$-dimensional embeddings by multiplying with an embedding matrix $\mathbf{E}\in\mathbb{R}^{d_{e}\times V}$, where the $i$-th column of $\mathbf{E}$ represents the embedding of the $i$-th symbol in $\mathcal{V}$. The resulting sequence of embeddings is denoted as $(\mathbf{e}_{1},\ldots,\mathbf{e}_{p})$, where $\mathbf{e}_{i}=\mathbf{E}\mathbf{x}_{i}\in\mathbb{R}^{d_{e}}$. The model processes the input embeddings using a mechanism called "self-attention". Each input embedding $\mathbf{e}_{i}$ is passed through a self-attention mechanism that calculates weighted relationships between all input embeddings – capturing the importance of each embedding relative to others. The model then aggregates these weighted representations to produce contextually enriched representations for each embedding. The contextually enriched representations produced by the self-attention mechanism are subsequently fed through feedforward neural networks (i.e., multilayer perceptrons, MLPs) to refine their information. Finally, the output tokens are generated based on the refined representations and converted back to human-readable format using the vocabulary $\mathcal{V}$.
@@ 3 Related Work
Interpreting and reverse engineering neural networks and transformers to find meaningful circuits has been an area of active research. Olah et al. [2020a] argued that by studying the connections between neurons and their weights, we can find meaningful algorithms (aka Circuits) in a “vision” neural network. Elhage et al. [2021] extended this approach to transformers, conceptualizing their operation in a mathematical framework that allows significant understanding of how transformer operate internally. Various tools [Foote et al., 2023, Conmy et al., 2023b, Garde et al., 2023] use this framework to semi-automate some aspects of reverse engineering. Nanda et al. [2023] reverse-engineered modular addition (e.g. 5 + 7 mod 10 = 2) showing the model used discrete Fourier transforms and trigonometric identities to convert modular addition to rotation about a circle.
@@
Figure 2: For 5-digit integer addition, these per-digit training loss curves show the model trains each answer digit semi-independently. Answer digit 0 (the “units” digit) is learnt much more quickly than other digits.
Circuit analysis can extract graphical circuit representations and analyze component interactions [Bau et al., 2017]. To enable analysis and interpretability, techniques in works like Petersen et al. [2021] symbolically reverse-engineer networks by recovering computational expressions. Research including Seth [2005] advocate analyzing networks causally, introducing structural causal models to infer mechanisms. Examinations of sequence models like Petersen et al. [2021] have analyzed the emergence and interaction of modular components during training. Lan and Barez [2024] analyzed and compared shared circuit subgraphs for related sequence continuation tasks in a large language model. Evolutionary perspectives such as Miikkulainen [2021] elucidate how selection pressures shape hierarchical representations. Lo et al. [2024] show how such representations, even if pruned, can quickly re-emerge in models after little re-training.
@@ 4 Methodology
A successful integer addition model must cope with a very large question and answer space. For 5 digit addition, there are 10 billion distinct questions (e.g. “54321+77779=") and 200,000 possible answers. Just one token (the “=") after seeing the complete question, the model must predict the first answer digit. It must correctly predict all 6 answers digits.
Figure 3: We refer to individual tokens in a 5-digit addition question as D4, .. D0, and D’4, .., D’0 and the answer tokens as A5, .., A0.
Our model was trained on 1.8 million out of 10 billion questions. After training, the model predicts answers to questions with low loss, showing the model does not rely on memorisation of training data. Fig. 2 shows the model trains each digit semi-independently suggesting the model performs integer addition by breaking down the task into parallel digit-specific streams of computation.
The Transformer model algorithms often differ significantly from our initial expectations. The easiest addition process for humans, given the digits can be processed in any order, is to add the unit digits first before moving on to higher value digits. This autoregressive transformer model processes text from left to right, so the model predicts the higher value digits (e.g. thousands) of the answer before the lower value digits (e.g. units). It can’t use the human process.
A key component of addition is the need to sum each digit in the first number with the corresponding digit in the second number. Transformer models contain “attention heads",the only computational sub-component of a model that can move information between positions (aka digits or tokens). Visualising which token(s) each attention head focused on in each row of the calculation provided insights. While our model works with 2, 3 or 4 attention heads, 3 attention heads give the most easily interpreted attention patterns. Fig. 4 shows the attention pattern for a single 5 digit addition calculation using 3 attention heads. Appendix C shows the same pattern for 10 and 15 digit addition. Appendix C shows the pattern with 2 or 4 attention heads.
@@ 5 Mathematical Framework
To help investigate, we created a mathematical framework describing what any algorithm must do if it is to perform addition correctly. Our intuition is that the model a) incrementally discovers a necessary and sufficient set of addition sub-tasks (minimising complexity), b) discovers these sub-tasks semi-independently (maximising parallelism), and c) treats each digit semi-independently (more parallelism). Our framework reflects this.
To explain the framework, Let $D=(d_{n-1},\ldots,d_{1},d_{0})$ and $D^{\prime}=(d^{\prime}_{n-1},\ldots,d^{\prime}_{1},d^{\prime}_{0})$ be two $n$-digitof digits. We assert that the framework utilizes three base functions that operate on individual digit pairs. The first is Base Add ( aka BA ), which calculates the sum of two digits $D_{i}$ and $D^{\prime}_{i}$ modulo 10, ignoring any carry over from previous columns. The second is Make Carry 1 ( aka MC1 ), which evaluates if adding digits $D_{i}$ and $D^{\prime}_{i}$ results in a carry over of 1 to the next column. The third is Make Sum 9 ( aka MS9 ), which evaluates if $D_{i}+D^{\prime}_{i}=9$ exactly.
@@
These tasks occur in the training data with different, predictable frequencies (e.g. BA is common, US9 is rarer). Compound tasks are reliant on the base tasks and so discovered later in training. The discovery of each task reduces the model loss by a different, predictable amount (e.g. BA by 50%, US9 by 5%). Combining these facts give an expected order of task discovery during training as shown in Fig. 5. We use this mathematical framework solely for analysis to gain insights. The model training and all loss calculations are completely independent of this mathematical framework.
@@ 6 Training Analysis
Fig. 2 shows the model trains each digit semi-independently. Armed with the mathematical framework, we investigated each digit separately. The Digit 0 calculation is the least interesting as it only uses BA (not UC1 or US9 ). Once discovered, Digit 0 always quickly refines to have the lowest loss and least noise (as expected). (Graphs in Appendix B.)
For the other digits, we categorised the training data into 3 non-overlapping subsets aligned to the BA, UC1 and US9 tasks, and graphed various combinations, finding interesting results. The US9 graphs are much noisier than other graphs (Fig. 6). We found that the model has low loss on simple US9 cases (e.g. 45 + 55 = 100) but has high loss on US9 cascades (e.g. 445 + 555 = 1000) where the MC1 must be propagated “right to left" two 3 or 4 columns. The model can’t perform these rare use cases safely, as it has a “left to right" algorithm.
@@
Both the BA and UC1 tasks need to move data between tokens, and so will be implemented in attention head(s). Fig. 7 shows they are trained semi-independently. We choose the number of attention heads in our model with the clearest separation of tasks in the attention pattern. We find (later) that our model has separate attention heads for the BA and UC1 tasks. Digit 4, the highest question digit, has a significantly different loss curve (shown in Fig. 8) than Digits 1, 2 and 3. This is partially explained by Digit 4 only having simple use cases (i.e. no US9 cascades). This does not explain the BA or UC1 differences. This difference persists with different seed values, and with 10 or 15 digit addition. We explain this difference later.
@@ 7 Prediction Analysis
Using the ablation interventions technique we overrode (mean ablated) the model memory (residual stream) at each row (aka token position) and confirmed that the addition algorithm does not use any data generated in rows 0 to 10 inclusive. In these rows the model has not yet seen the full question and every digit in the question is independent of every other digit, making accurate answer prediction infeasible. The model also does not use the last (17th) row. Therefore, the addition answer calculation is started and completed in 6 rows (11 to 16). Further ablation experiments confirmed that the A0 to A4 answers are calculated one row before being revealed. (Details in Appendix H.)
The model has slightly different algorithms for the first answer digits (A5 and A4), the middle answer digits (A3 and A2) and the last answer digits (A1 and A0). Fig. 1 has a simplified version of how the model calculates the answer digit A3. Fig. 9 has more details. For 5 digit addition, there are 2 middle answer digits (A3 and A2) whereas for 15 digit addition there are 12 middle answer digits that use this algorithm.
Figure 9: A: To predict answer digit A3, the addition algorithm must combine information from digits 3, 2 and 1. B: The 1st head calculates MC1 on digit 1. C: The 2nd head calculates MC1 and MS9 (at most one of which can be true) on digit 2. D: The 3rd head calculates Base Add on digit 3. E: The MLP layer uses trigrams to combine the information from the 3 heads to give the final answer A3, one row before it is output. Appendix G shows this algorithm as pseudocode.
The A3 addition algorithm has three clauses related to digits 3, 2 and 1. Ablating each head in turn shows that the 3rd head has most impact on loss, the 2nd head has less impact, and the 1st head has little impact. This aligns with the intuition that the sum “D3 + D’3" matters most, the MC1 from the previous digit (D2 + D’2) matters less, and the rare MC1 from the previous previous digit (D1 + D’1) matters least. The last two answer digits, A1 and A0, use a simplified a version of the A3 algorithm, as some clauses are not necessary.
The A3 algorithm can also be applied to A4. But the Digit 4 training curve is better (faster) than the middle digits. The attention patterns show that for A4, the model is using all the heads in row 11 (the “=" input token) when the A3 algorithm doesn’t require this. Uniquely, A4 utilises more “compute" than is available to A3, A2, A1 or A0. We assume the model uses this advantage to implement a faster-training and lower-loss algorithm for A5 and A4. We haven’t worked out the details of this.
@@
Despite being feasible, the model does not calculate the task MC1 in rows 7 to 11. Instead, it completes each answer digit calculation in 1 row, possibly because there are training optimisation benefits in generating a “compact" algorithm.
@@
8 Algorithm Reuse
We explored whether the above algorithm is learned by similar models. We trained a separate 1-layer model for 5-digit addition with a different random seed and optimization algorithm. The new model’s answer format requires it to predict "+" as the first answer token, e.g., "12345+67890=+080235". Despite these changes, Figures 24,  25, and  26 show that the new model’s behavior has many similarities to the previous model.
Intervention ablation demonstrates that the new model uses the $\operatorname{BaseAdd}$, $\operatorname{MakeCarry}$, and $\operatorname{MakeSum}$ sub-tasks in the same way as the previous model. The new model exhibits the same strengths and weaknesses, e.g., it can calculate a simple $\operatorname{MS}$ case but not a cascading $\operatorname{MS}$ case. We claim that the new and previous models use essentially the same algorithm, with minor variations.
Our analysis suggests that the transformer architecture, when trained on the addition task, converges to a consistent algorithm for performing digit-wise addition. This algorithm leverages the self-attention mechanism to discover and execute the necessary sub-tasks in a parallel and semi-independent manner. Despite differences in random initialization, optimization algorithms, and output formatting, the models exhibit similar internal behavior and capabilities, indicating a robust algorithmic solution emerges from the transformer’s architecture and training process.
9 Conclusions